### v1

@@ C.3 Final NEGFC results
Table 2 shows that the astrometry and photometry extractions using NEGFC+roll and NEGFC+IPCA are consistent in all cases where retrieval of planet parameters could be made in both the roll-subtracted and IPCA images (i.e., planet with both filters, and planet in the F480M data). This suggests that IPCA recovered most of the self- and over-subtracted flux affecting roll-subtracted images. Nonetheless, neither approach is devoid of weakness for parameter estimation. It is unclear whether IPCA did recover all the self-/over-subtracted flux for cirumstellar signals, including the protoplanets, while on the other hand, it is unclear how much residual extended signals filtered by roll subtraction may lead to a mis-estimation of the NEGFC+roll planet fluxes. We therefore conservatively consider as final results (presented in Table 1) the mean of the results obtained with NEGFC+roll and NEGFC+IPCA, in cases where both estimates are available, and adopt the largest uncertainties among the two approaches.
@@
In general, the visual vetting of NEGFC+IPCA results rejected the results obtained with the sum figure of merit as it systematically led to a ‘hole’ within the surrounding patch of pixels, which we interpret as likely overestimating the flux of the protoplanet alone. Only in the case of planet in the F187N data, did the results appear visually satisfactory as the protoplanet is located just outward of the inner disk (although signals from both components appear connected at the resolution of our observations). In that case, it was fair to consider the possibilities that either most of the flux at the location of within our flux uncertainties is due to the protoplanet itself, or that there is non-zero contributing signal from the inner disk. These possibilities are well captured with the sum and standard deviation figures of merit, respectively. We therefore considered the median and standard deviation of the parameters retrieved for all corresponding subcases as our reported NEGFC+IPCA results. We also note that this case (planet in F187N) was the only one where the Hessian-based figure of merit did not work properly for NEGFC+IPCA, due to similar flux levels for and the adjacent inner disk signals. On the contrary, this figure of merit worked the best in the NEGFC+roll approach, compared to sum or standard deviation figures of merit, based on our positive fake companion injection tests, which we interpret as due to the stronger self-subtraction of (the closer-in) inner disk signals, which makes the planet signal stand out from it. Similar remarks can be made for planet in the F480M images, where the IPCA images better recover inner disk signals, hence tend to provide slightly closer radial separation estimates for the protoplanet due to the bias from the inner disk.
@@
For planet in the F187N data, the estimated parameters are affected by large uncertainties as a lot of additional signals is recovered by IPCA around the planet, including the spiral accretion stream and candidate , which makes the estimate of the contribution from the planet alone difficult. For planet in the F480M data, the NEGFC+roll and NEGFC+IPCA results are very similar, both approaches leading to visually satisfactory results in the images obtained after subtraction of the respective optimal planet parameters. We note nevertheless a potential shift to a larger estimated PA in the F480M data than expected from orbital fits, which may reflect either contamination by the spiral accretion stream in this lower angular resolution image, or a mis-estimation of the orbit from earlier astrometric measurements.
@@ C.4 Additional tests
Regarding the presence of excess signals compared to atmospheric model predictions for the protoplanets (Sec. 3.2), we subtracted a scaled version of the observed PSF with the appropriate contrast ratio to match the flux predicted by the best-fit DRIFT-PHOENIX model for the SED of planet presented in Wang et al. (2021b), at the predicted location for the planet based on the orbital fits presented in Wang et al. (2021b). Using the position inferred by NEGFC in the previous sections, instead of the predicted location, does not change the conclusion. The results of this safety check are shown in the left and right columns of Fig. 9 for PDS 70 c, in the F187N and F480M filter images, respectively. The residual signals observed at the location of the protoplanet suggest indeed the presence of excess emission compared to predictions from the best-fit atmospheric models alone.
We performed a similar test for planet in the F480M data, shown in the middle column of Fig. 9. In this case, we subtracted the flux corresponding to the lower uncertainty of the estimate found with NEGFC (reported in Table 1) which is roughly midway between the best-fit BT-SETTL model with blackbody excess, and both the DRIFT-PHOENIX and extinct EXOREM models without blackbody excess. This test suggests that our reported uncertainty is sound, and that there is thus tentative excess compared to predictions from atmospheric models without an extra blackbody contribution representative of circumplanetary disk emission.
For comparison, the images obtained after subtraction of the optimal parameters found by NEGFC for protoplanets and are shown in Fig. 1c and f for the F187N and F480M images, respectively.
@@ Appendix D Predicted and observed spirals
Figure 10 compares our disk-subtracted F187N and F480M images (panels b and c) to a simulated IR image predicting the spiral accretion stream associated with protoplanet (panel a). The latter is based on dedicated 3D hydro-dynamical simulations made with smoothed-particle hydrodynamics code PHANTOM (Price et al. 2018), presented in Toci et al. (2020). The simulated image corresponds to a radiative-transfer prediction at 2.11 m of the PHANTOM simulation made with Monte-Carlo radiative transfer code MCFOST (Pinte et al. 2006), where a proxy for the outer disk signal has been subtracted. This proxy consists in another radiative-transfer prediction for a different snapshot of the same hydro-dynamical simulation corresponding to planet being located at 180 from its observed position angle. We therefore mask the bottom part of the image, as it is affected by strong residuals associated to the spiral accretion stream being subtracted from that part of the image. We measured the trace of the spiral accretion stream in the simulated image, through the identification of local radial maxima in the image, in steps of 1 . The trace, shown with white dots, is then also plotted on top of the F187N and F480M images. We notice a remarkable correspondence with the observed spiral-like signal in the F187N image, and with the inner arm of the tentative fork seen in the F480M image (better seen in Fig. 1e and f)
The arm-like signal identified and characterized in Juillard et al. (2022) is also shown with orange dots in Fig. 10c. This corresponds to the trace inferred from a 2021 H-band VLT/SPHERE dataset (ESO Program 60.A-9801; see more details in Juillard et al. 2022). We also note a good agreement with the outer arm observed in the F480M image, suggesting it traces the same feature as observed in VLT/SPHERE images. The coarse angular resolution of the image and the different wavelength of the observation (compared to prior ground-based images where the feature is seen) prevent however a meaningful proper motion analysis of the spiral feature similar to what was done in Juillard et al. (2022).
@@ E.1 Signal-to-Noise Ratio maps
Figure 11 shows the Signal-to-Noise Ratio (SNR) maps obtained with roll subtraction on the image cube where the optimal radiative disk model, and protoplanets and c, were all subtracted with their optimal parameters found with NEGFD (Appendix B) and NEGFC (Appendix C), respectively. Since both protoplanets share similar projected radii, the SNR maps are shown with planet (resp. ) removed to evaluate the significance of planet (resp. ) in the left (resp. middle) panel. Both planets are removed in the right panel, to assess the SNR of residual signals apart from the planets. We do not detect any additional planet candidate in the outer disk. Only the spiral-like feature located in the outer wake of planet stands out at SNR.
@@ E.2 Mass sensitivity curves
Fig. 12 (left) shows the contrast curves achieved with PCA-roll (Appendix A.2) and TRAP (Samland et al. 2021) on the image cube obtained after subtraction of the optimal disk and planet models, determined with NEGFD and NEGFC, respectively. A wedge was defined to encompass PA values between 10 and 160 when calculating the contrast curves, to avoid any significant disk residuals, including the arm-like feature, to bias the contrast estimation.
@@
We converted the achieved contrasts to mass sensitivities using the ATMO2020 models (Phillips et al. 2020), with both equilibrium and non-equilibrium chemistry, and the BEX models (Linder et al. 2019), assuming an age of 5.41 Myr (Müller et al. 2018). These mass limits are shown in Fig. 12 (right), with the shaded region representing the 1 uncertainty on mass limits due to the uncertainty in target age. The ATMO2020 evolutionary models cover masses above 0.524, while the BEX models cover masses below 2. There is some scatter in the calculated mass sensitivity based on the different spectra for each model assumption; models at this young age are also finely sensitive to initial conditions. The curves show that we would have been sensitive to planets with a mass (resp. 0.8) at 0.6 (resp. 1) separation from the star, according to the BEX and ATMO (in chemical equilibrium) models – noting that we hit the bottom of the ATMO grid of models (in chemical equilibrium) in terms of planet mass. In case of non-equilibrium chemistry in the atmosphere, planets as massive as 2-4 may still be present, yet unseen, in the outer disk. Even larger mass giant planets may be hidden in the outer disk if their signal is extinct. Nonetheless, multi-Jovian mass giant planets would likely also have affected the gas density profile, which ALMA reveals to extend beyond (i.e., beyond 170 au) in radius without any additional gaps than the one carved by protoplanets and (Keppler et al. 2019).
Appendix F The F187N excess for planet c
@@
All flux estimates shown in Figs. 2 and 3, but the NIRCam ones rely on either a model of the star plus inner disk or absolute (spectro-)photometry of the star measured at a different epoch than the epoch of the observation which allowed the contrast measurement of the planet with respect to the star, for the scaling of contrasts to absolute fluxes. Variability of the star plus unresolved inner disk flux (Casassus & Cárcamo 2022; Perotti et al. 2023) can thus result in significant differences between the actual photometry and the assumed one for the star, hence affect accordingly the protoplanet flux. To illustrate the amplitude of this bias, we multiplied our measured contrasts for and to the F187N stellar flux inferred based on the 0.7-2.7m SpeX spectrum of PDS 70 (Long et al. 2018) used in other works, and show the results with light blue error bars in Figs. 2 and 3. This may thus partly account for the observed discrepancy.
@@
Another source of misestimation for the flux of planet is the bright outer disk edge, for measurements leveraging angular differential imaging (Marois et al. 2006), and done without prior disk model subtraction from the images. Bright disk signals captured in the model stellar PSF lead to a pair of negative traces encasing the positive trace of the disk in PSF-subtracted images, similar to what can be seen in the roll-subtraction images (Fig. 4a and e), significantly dampening any other signal in the negative trace (e.g., planet and the accretion stream in Fig. 4a). Classical NEGFC (with the ‘sum’ figure of merit) or other point-source forward-modeling approaches are not tailored to properly account for self- and over-subtraction effects associated to the bright disk, and may underestimate the point source flux if located in such very negative disk trace. The exact amplitude of this bias is difficult to assess as it depends on the disk scattering phase function (stronger effect close to PA where the disk is the brightest), the number of principal components used for PCA-based reductions (see e.g., Fig. 2 in Juillard et al. 2022), and the wavelength of the observation, as the disk is brighter, hence the effect stronger, at shorter wavelengths. If the SPHERE/IFS measurements (0.9–1.6 m) are underestimated, the best-fit models presented in Wang et al. (2021b) may also underpredict the F187N flux, hence inflate our reported 1.87m excess.
@@
Another possibility is the inclusion of bright circumplanetary signals that are resolved in ground-based observations, but unresolved in the lower angular resolution images of our observations, hence contributing only to our NIRCam measurements. However, the difference in angular resolution between the longest wavelengths of the SPHERE/IFS observations and our JWST/NIRCam images is minor, with an estimated FWHM about 70% smaller for the former compared to the latter. This hypothesis appears thus unlikely to account on its own for the factor excess measured in the F187N compared to atmospheric model predictions.
@@
Alternatively, part of the measured F187N excess may be assigned to the presence of significant Pa- line emission, an hypothesis discussed in more details in the main text (Sec. 3.2).
@@ Appendix G Schematic representation of the system
In Figure 13, we summarize our knowledge about and interpretation of the different features of the PDS 70 system, drawing on Keppler et al. (2018, 2019), Mesa et al. (2019), Isella et al. (2019), Benisty et al. (2021), Casassus & Cárcamo (2022), Wang et al. (2021b), and this work.

### v2

@@ C.3 Final NEGFC results
Table 2 shows that the astrometry and photometry extractions using NEGFC+roll and NEGFC+IPCA are consistent in all cases when retrieval of planet parameters could be made in both the roll-subtracted and IPCA images (i.e., planet with both filters, and planet in the F480M data). This suggests that IPCA recovered most of the self- and oversubtracted flux affecting roll-subtracted images. Nonetheless, neither approach is devoid of weakness for the parameter estimation. It is unclear whether IPCA did recover all the self- or oversubtracted flux for cirumstellar signals, including the protoplanets, while on the other hand, it is unclear how much residual extended signals filtered by roll subtraction may lead to a misestimation of the NEGFC+roll planet fluxes. We therefore conservatively consider (presented in Table 1) the mean of the results obtained with NEGFC+roll and NEGFC+IPCA as our final results when both estimates are available, and we adopt the larger uncertainties of the two approaches.
@@
In general, the visual vetting of NEGFC+IPCA results rejected the results obtained with the sum figure of merit as it systematically led to a hole within the surrounding patch of pixels, which we interpret as likely overestimating the flux of the protoplanet alone. Only in the case of planet in the F187N data did the results appear visually satisfactory because the protoplanet is located just outward of the inner disk (although the signals from the two components appear to be connected at the resolution of our observations). In this case, it was fair to consider the possibilities that either most of the flux at the location of within our flux uncertainties is due to the protoplanet itself, or that there is nonzero contributing signal from the inner disk. These possibilities are well captured with the sum and standard deviation figures of merit, respectively. We therefore considered the median and standard deviation of the parameters retrieved for all corresponding subcases as our reported NEGFC+IPCA results. We also note that in this case (planet in F187N) alone, the Hessian-based figure of merit did not work properly for NEGFC+IPCA because the flux levels for and the adjacent inner-disk signals were similar. In contrast, this figure of merit worked the best in the NEGFC+roll approach compared to the sum or standard deviation figures of merit, based on our positive fake companion injection tests, which we interpret as due to the stronger self-subtraction of (the closer-in) inner-disk signals, which causes the planet signal to stand out from it. Similar remarks can be made for planet in the F480M images, where the IPCA images recover inner-disk signals better and therefore tend to provide slightly closer radial separation estimates for the protoplanet due to the bias from the inner disk.
@@
For planet in the F187N data, the estimated parameters are affected by large uncertainties because many additional signals are recovered by IPCA around the planet, including the spiral accretion stream and candidate , which makes the estimate of the contribution from the planet alone difficult. For planet in the F480M data, the NEGFC+roll and NEGFC+IPCA results are very similar. Both approaches lead to visually satisfactory results in the images obtained after subtracting the respective optimal planet parameters. We nevertheless note a potential shift to a larger estimated PA in the F480M data than expected from orbital fits, which may either reflect contamination by the spiral accretion stream in this lower angular resolution image or a misestimation of the orbit from earlier astrometric measurements.
@@ C.4 Additional tests
To determine excess signals compared to atmospheric model predictions for the protoplanets (Sec. 3.2), we subtracted a scaled version of the observed PSF with the appropriate contrast ratio to match the flux predicted by the best-fit Drift-Phoenix model for the SED of planet presented in Wang et al. (2021b) at the predicted location for the planet based on the orbital fits presented in Wang et al. (2021b). The position inferred by NEGFC in the previous sections instead of the predicted location does not change the conclusion. The results of this safety check are shown in the left and right columns of Fig. 9 for PDS 70 c in the F187N and F480M filter images, respectively. The residual signals observed at the location of the protoplanet indeed suggest excess emission compared to predictions from the best-fit atmospheric models alone.
We performed a similar test for planet in the F480M data, shown in the middle column of Fig. 9. In this case, we subtracted the flux corresponding to the lower uncertainty of the estimate found with NEGFC (reported in Table 1), which is roughly midway between the best-fit BT-Settl model with blackbody excess and both the Drift-Phoenix and extinct EXOREM models without blackbody excess. This test suggests that our reported uncertainty is sound, and that there is thus tentative excess compared to predictions from atmospheric models without an additional blackbody contribution representative of circumplanetary disk emission.
For comparison, the F187N and F480M images obtained after subtracting protoplanets and with the optimal parameters found by NEGFC are shown in Fig. 1c and f.
@@ Appendix D Predicted and observed spirals
Figure 10 compares our disk-subtracted F187N and F480M images (panels b and c) to a simulated IR image predicting the spiral accretion stream associated with protoplanet (panel a). The latter is based on dedicated 3D hydrodynamical simulations made with the smoothed-particle hydrodynamics code PHANTOM (Price et al. 2018), presented in Toci et al. (2020). The simulated image corresponds to a radiative transfer prediction at 2.11 m of the PHANTOM simulation made with the Monte Carlo radiative transfer code MCFOST (Pinte et al. 2006), where a proxy for the outer-disk signal was subtracted. This proxy consists of another radiative transfer prediction for a different snapshot of the same hydrodynamical simulation, corresponding to planet being located at 180 from its observed position angle. We therefore masked the bottom part of the image because it is affected by strong residuals associated with the spiral accretion stream that was subtracted from that part of the image. We measured the trace of the spiral accretion stream in the simulated image by identifying local radial maxima in the image in steps of 1 . The trace, shown with white dots, is then also plotted on top of the F187N and F480M images. The correspondence with the observed spiral-like signal in the F187N image and with the inner arm of the tentative fork seen in the F480M image is remarkable (better seen in Fig. 1e and f)
The arm-like signal identified and characterized in Juillard et al. (2022) is also shown with orange dots in Fig. 10c. This corresponds to the trace inferred from a 2021 H-band VLT/SPHERE dataset (ESO Program 60.A-9801; see more details in Juillard et al. 2022). We also note a good agreement with the outer arm observed in the F480M image, suggesting that it traces the same feature as observed in VLT/SPHERE images. The coarse angular resolution of the image and the different wavelength of the observation (compared to prior ground-based images in which the feature is seen) prevents a meaningful proper motion analysis of the spiral feature similar to what was done in Juillard et al. (2022), however.
@@ E.1 Signal-to-noise ratio maps
Figure 11 shows the S/N maps obtained with roll subtraction on the image cube from which the optimal radiative disk model and protoplanets and c were all subtracted with their optimal parameters found with NEGFD (Appendix B) and NEGFC (Appendix C), respectively. Because the protoplanets share similar projected radii, the S/N maps are shown without planet () to evaluate the significance of planet () in the left (middle) panel. Both planets are removed in the right panel to assess the S/N of the residual signals apart from the planets. We do not detect any additional planet candidate in the outer disk. Only the spiral-like feature located in the outer wake of planet stands out at S/N.
@@ E.2 Mass sensitivity curves
Fig. 12 (left) shows the contrast curves achieved with PCA-roll (Appendix A.2) and TRAP (Samland et al. 2021) on the image cube obtained after subtracting the optimal disk and planet models determined with NEGFD and NEGFC, respectively. A wedge was defined to encompass PA values between 10 and 160 when calculating the contrast curves to avoid any significant disk residuals, including the arm-like feature, from biasing the contrast estimation.
@@
We converted the achieved contrasts into mass sensitivities using the ATMO2020 models (Phillips et al. 2020) with both equilibrium and nonequilibrium chemistry, and the BEX models (Linder et al. 2019) assuming an age of 5.41 Myr (Müller et al. 2018). These mass limits are shown in Fig. 12 (right), in which the shaded region represents the 1 uncertainty on the mass limits due to the uncertainty in target age. The ATMO2020 evolutionary models cover masses above 0.524, and the BEX models cover masses below 2. There is some scatter in the calculated mass sensitivity based on the different spectra for each model assumption; models at this young age are also highly sensitive to initial conditions. The curves show that we would have been sensitive to planets with a mass (0.8) at 0.6 (resp. 1) separation from the star, according to the BEX and ATMO (in chemical equilibrium) models; we note that we reached the bottom of the ATMO grid of models (in chemical equilibrium) in terms of planet mass. In case of nonequilibrium chemistry in the atmosphere, planets as massive as 2-4 may still be present, if unseen, in the outer disk. Even higher-mass giant planets may be hidden in the outer disk if their signal is extinct. Nonetheless, multi-Jovian mass giant planets would likely also have affected the gas density profile, which ALMA reveals to extend beyond (i.e., beyond 170 au) in radius without any additional gaps than the gap that is carved by protoplanets and (Keppler et al. 2019).
Appendix F F187N excess for planet c
@@
All flux estimates shown in Figs. 2 and 3 except for those from the NIRCam rely on either a model of the star plus inner disk or on absolute (spectro-)photometry of the star measured at a different epoch than the epoch of the observation that allowed the contrast measurement of the planet with respect to the star for the scaling of contrasts to absolute fluxes. The stellar variability plus unresolved inner-disk flux (Casassus & Cárcamo 2022; Perotti et al. 2023) can thus result in significant differences between the actual photometry and the assumed photometry for the star, and may therefore accordingly affect the protoplanet flux. To illustrate the amplitude of this bias, we multiplied our measured contrasts for and by the F187N stellar flux inferred based on the 0.7-2.7m SpeX spectrum of PDS 70 (Long et al. 2018) used in other works. We show the results with light blue error bars in Figs. 2 and 3. This may thus partly account for the observed discrepancy.
@@
Another source of misestimation for the flux of planet is the bright outer-disk edge for measurements leveraging angular differential imaging (Marois et al. 2006), and performed without prior disk model subtraction from the images. Bright disk signals captured in the model stellar PSF lead to a pair of negative traces encasing the positive trace of the disk in PSF-subtracted images, similar to what can be seen in the roll-subtraction images (Fig. 4a and e). This significantly dampens any other signal in the negative trace (e.g., planet and the accretion stream in Fig. 4a). Classical NEGFC (with the sum figure of merit) or other point-source forward-modeling approaches are not tailored to properly account for self- and oversubtraction effects associated with the bright disk and may underestimate the point-source flux when it is located in a very negative disk trace like this. The exact amplitude of this bias is difficult to assess as it depends on the disk scattering phase function (stronger effect close to PAb where the disk is brightest), the number of principal components used for PCA-based reductions (see e.g., Fig. 2 in Juillard et al. 2022), and the wavelength of the observation because the disk is brighter, hence the effect stronger, at shorter wavelengths. If the SPHERE/IFS measurements (0.9–1.6 m) are underestimated, the best-fit models presented in Wang et al. (2021b) may also underpredict the F187N flux, hence inflate our reported 1.87m excess.
@@
Another possibility is the inclusion of bright circumplanetary signals that are resolved in ground-based observations, but are unresolved in the lower angular resolution images of our observations, hence contributing only to our NIRCam measurements. However, the difference in angular resolution between the longest wavelengths of the SPHERE/IFS observations and our JWST/NIRCam images is minor, with an estimated FWHM about 70% smaller for the former compared to the latter. This hypothesis therefore appears to be unlikely to account alone for the excess of a factor of measured in the F187N compared to atmospheric model predictions.
@@
Alternatively, part of the measured F187N excess may be assigned to the presence of significant Pa- line emission. This hypothesis is discussed in more details in the main text (Sec. 3.2).